# An explicit local geometric Langlands for supercuspidal representations: the toral case

### v1

@@ 6.3. Central characters
In the current situation, $J^{+}$ is just the kernel of the projection from $J$ to $S(1)/S(N+1)$. Explicitly, its Lie algebra is
$\mathfrak{j}^{+}=\begin{cases}\bigoplus_{i=\frac{N+1}{2}}^{N}\tau_{i}u^{i}\oplus\mathfrak{p}(N+1),\hskip 49.79231ptN\text{ odd,}\\ \bigoplus_{i=\frac{N}{2}+1}^{N}\tau_{i}u^{i}\oplus\mathfrak{m}u^{\frac{N}{2}}\oplus\mathfrak{p}(N+1),\quad N\text{ even.}\end{cases}$
By construction, $\tilde{\phi}(J^{+})=1$. Recall from (31) that $\bar{\mathfrak{j}}:=\mathfrak{j}/\mathfrak{j}^{+}\simeq\mathfrak{s}(1)/\mathfrak{s}(N+1)\simeq\bigoplus_{1\leq i\leq N}\mathfrak{t}_{Y,i}u^{i}$.
@@ 7.2. The $K$-type of the central block
To sum up, for toral supercuspidal representations and irreducible isoclinic connections, Conjecture.(i)-(iii) 13 has been established in Lemma 33 and Theorem 34. The part (iv) of the conjecture follows from the same proof as Lemma 33.(ii).

### v2

@@ 6.3. Central characters
Recall we defined a subspace $\mathfrak{j}^{\prime}\subset\mathfrak{j}$ (30). Explicitly, in the current situation it is given by
$\mathfrak{j}^{\prime}=\begin{cases}\bigoplus_{i=\frac{N+1}{2}}^{N}\tau_{i}u^{i}\oplus\mathfrak{p}(N+1),\hskip 49.79231ptN\text{ odd,}\\ \bigoplus_{i=\frac{N}{2}+1}^{N}\tau_{i}u^{i}\oplus\mathfrak{m}u^{\frac{N}{2}}\oplus\mathfrak{p}(N+1),\quad N\text{ even.}\end{cases}$
Lemma 28.
The subspace $\mathfrak{j}^{\prime}\subset\mathfrak{j}$ is an ideal, so that $\mathfrak{j}^{\prime}=\mathfrak{j}^{+}$.
Proof.
We need to show $[\mathfrak{j},\mathfrak{j}^{\prime}]\subset\mathfrak{j}^{\prime}$. Observe $\mathfrak{j}=\mathfrak{j}^{\prime}+\mathfrak{s}(1)$. It is easy to see that the adjoint action of $\mathfrak{s}(1)$ preserves $\mathfrak{j}^{\prime}$. To show $[\mathfrak{j}^{\prime},\mathfrak{j}^{\prime}]\subset\mathfrak{j}^{\prime}$, from the depth consideration it suffices to show $[\mathfrak{m},\mathfrak{m}]\subset\tau_{N}$ when $N$ is even. This follows from the construction of $\mathfrak{m}$ (50). ∎
Therefore, the quotient (31) becomes $\bar{\mathfrak{j}}=\mathfrak{j}/\mathfrak{j}^{+}\simeq\mathfrak{s}(1)/\mathfrak{s}(N+1)\simeq\bigoplus_{1\leq i\leq N}\mathfrak{t}_{Y,i}u^{i}$.
@@ 7.2. The $K$-type of the central block
To sum up, for toral supercuspidal representations and irreducible isoclinic connections, Conjecture.(i)-(iii) 14 has been established in Lemma 35 and Theorem 36. The part (iv) of the conjecture follows from the same proof as Lemma 35.(ii).